# EG-SpikeFormer: Eye-Gaze Guided Transformer on Spiking Neural Networks for Medical Image Analysis

### v1

@@ Abstract
Neuromorphic computing has emerged as a promising energy-efficient alternative to traditional artificial intelligence, predominantly utilizing spiking neural networks (SNNs) implemented on neuromorphic hardware. Significant advancements have been made in SNN-based convolutional neural networks (CNNs) and Transformer architectures. However, their applications in the medical imaging domain remain underexplored. In this study, we introduce EG-SpikeFormer, an SNN architecture designed for clinical tasks that integrates eye-gaze data to guide the model’s focus on diagnostically relevant regions in medical images. This approach effectively addresses shortcut learning issues commonly observed in conventional models, especially in scenarios with limited clinical data and high demands for model reliability, generalizability, and transparency. Our EG-SpikeFormer not only demonstrates superior energy efficiency and performance in medical image classification tasks but also enhances clinical relevance. By incorporating eye-gaze data, the model improves interpretability and generalization, opening new directions for the application of neuromorphic computing in healthcare.
@@ 1 Introduction
In this paper, we propose a novel gaze-guided spike-driven hybrid model EG-SpikeFormer for medical diagnosis for the first time, which integrates the low-power computation benefits of SNNs with the powerful feature extraction capabilities of Transformers. The model incorporates radiologists’ eye-gaze data as prior information during training, effectively guiding the model’s attention. Our experiments on two public medical datasets demonstrate the model’s superior performance in both energy efficiency and diagnostic accuracy. The main contributions of our work are as follows:
@@
  • To the best of our knowledge, this is the first attempt to introduce a hybrid model combining SNN and Transformer in the field of medical diagnosis, leveraging both low power consumption and strong interpretability.
@@ 2.1 Blocks with Spike Neurons
The Transformer-based SNN block, depicted in Figure 3, integrates the attention mechanisms of traditional Transformers with the temporal dynamics of spiking neurons. In this block, spiking neurons generate spike-based queries $Q_{s}$, keys $K_{s}$, and values $V_{s}$, which are essential components of the self-attention mechanism. These spike-based representations are obtained by processing incoming spike signals through re-parameterized convolution layers that culminate with spiking neurons. Specially, for a input $X$ to the transformer-based SNN block, the queries, keys, and values are computed as
@@ 2.2 Eye-gaze Guided Spike Vision Transformer
We augment the original images by integrating eye-gaze data to highlight key regions without losing information from other areas. Let $\mathbf{I}\in\mathbb{R}^{H\times W\times C}$ denote the original image, and let $\mathbf{M}\in\mathbb{R}^{H\times W}$ be the normalized eye-gaze mask with values between 0 and 1. The enhanced image $\mathbf{I}^{\prime}$ is computed as:
@@
where $\odot$ denotes element-wise multiplication, $\mathbf{1}$ is an all-ones matrix of compatible dimensions, and $\alpha$ is a hyperparameter controlling the emphasis on eye-gaze regions. This operation amplifies pixel intensities in areas with higher eye-gaze values, effectively guiding the network’s focus.
To align the model’s attention with human gaze patterns, we introduce an attention alignment loss. Let $\mathbf{A}_{t}\in\mathbb{R}^{N\times N}$ be the model’s attention map from the last layer of the final Transformer block, and $\mathbf{A}_{g}\in\mathbb{R}^{N\times N}$ be the attention map derived from eye-gaze data, where $N$ is the number of tokens. The attention alignment loss $\mathcal{L}_{\text{align}}$ is defined as:
@@ 3.2 Results and Analysis
As shown in Table 1, our eye-gaze guided SNN architecture, EG-SpikeFormer achieves state-of-the-art performance across all metrics, outperforming classical ViTs, eye-gaze augmented ViTs, and CNNs. Despite lower energy consumption, our model enhances performance, demonstrating efficiency and superior capability in complex medical imaging tasks. Leveraging the spatio-temporal event-driven computation of SNNs, our architecture balances high performance with low energy consumption without compromising accuracy. It surpasses traditional deep learning architectures, highlighting the potential of SNNs in resource-constrained medical environments. Integrating eye-gaze data effectively guides the model’s attention, improving accuracy and convergence speed without significantly increasing computational resources. We evaluated two training processes: a two-step and a four-step. Even with two-step training, our model matches or exceeds other state-of-the-art methods. The four-step training further enhances capabilities, indicating that additional steps can unlock more potential without substantially increasing energy consumption.
@@ 3.3 Ablation Study
We conducted an ablation study to assess the impact of the Gaze Mask (GM), Attention Mapping (AM), and the four-step training process (T4) on model performance. The detailed results are presented in Table 2. The model without T4 also showed lower performance metrics. The highest performance was achieved when all components were included, confirming that GM, AM, and T4 each contribute significantly to the model’s effectiveness.
@@
GM AM T4 Acc. $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ Acc. $\uparrow$ AUC $\uparrow$ F1 $\uparrow$
@@ 4 Discussion
Neuromorphic chips, leveraging SNNs as the most accessible approach, offer a transformative solution for clinical diagnostics, where energy efficiency, data scarcity, and interpretability are crucial [25, 26]. EG-SpikeFormer excels in overcoming these challenges by addressing shortcut learning and ensuring focus on diagnostically relevant features, making it highly suitable for real-time medical applications in resource-constrained environments. More critically, EG-SpikeFormer offers a forward-thinking framework for neuromorphic chip design. By leveraging eye-gaze-driven, spatio-temporal event processing, it shows how neuromorphic systems can be optimized for clinical tasks. This model not only enhances diagnostic accuracy but also drives a targeted approach to algorithm-hardware co-design, enabling the creation of chips that are both computationally efficient and highly effective in medical settings. Furthermore, the model’s ability to reduce data requirements and computational overhead offers significant cost and time savings in chip development. By guiding the design of energy-efficient and scalable neuromorphic architectures, EG-SpikeFormer accelerates the path toward more accessible, high-performance clinical diagnostics, ultimately expanding the potential of neuromorphic technologies in healthcare.
@@ 5 Conclusion
EG-SpikeFormer integrates eye-gaze data with SNNs to enhance medical image analysis, leveraging SNNs’ energy efficiency and temporal dynamics with Transformers’ feature extraction capabilities. This model achieves superior diagnostic accuracy and interpretability by directing attention to clinically relevant regions, improving generalizability and addressing the limitations of conventional models. These results highlight neuromorphic computing’s potential in healthcare, providing a high-performance, low-energy solution, particularly suited for data-constrained medical environments.

### v2

@@ Abstract
Neuromorphic computing has emerged as a promising energy-efficient alternative to traditional artificial intelligence, predominantly utilizing spiking neural networks (SNNs) implemented on neuromorphic hardware. Significant advancements have been made in SNN-based convolutional neural networks (CNNs) and Transformer architectures. However, neuromorphic computing for the medical imaging domain remains underexplored. In this study, we introduce EG-SpikeFormer, an SNN architecture tailored for clinical tasks that incorporates eye-gaze data to guide the model’s attention to the diagnostically relevant regions in medical images. Our developed approach effectively addresses shortcut learning issues commonly observed in conventional models, especially in scenarios with limited clinical data and high demands for model reliability, generalizability, and transparency. Our EG-SpikeFormer not only demonstrates superior energy efficiency and performance in medical image prediction tasks but also enhances clinical relevance through multi-modal information alignment. By incorporating eye-gaze data, the model improves interpretability and generalization, opening new directions for applying neuromorphic computing in healthcare.
@@ 1 Introduction
In this paper, we propose a novel gaze-guided spike-driven hybrid model EG-SpikeFormer for medical diagnosis for the first time, which integrates the low-power computation benefits of SNN with the powerful feature extraction capabilities of Transformers. The model incorporates radiologists’ eye-gaze data as prior information during training, effectively guiding the model’s attention. Our experiments on two public medical datasets demonstrate the model’s superior performance in both energy efficiency and diagnostic accuracy. The main contributions of our work are as follows:
@@
  • To the best of our knowledge, this is the first attempt to introduce a hybrid SNN model combining CNN and Transformer in the field of medical diagnosis, leveraging both low power consumption and strong interpretability.
@@
  • We firstly propose a hardware-aware co-design framework that integrates neuromorphic computing and eye-gaze guidance, addressing medical challenges such as shortcut learning and data scarcity. This collaborative design not only enhances diagnostic accuracy and energy efficiency but also underscores the significance of such synergies in advancing practical healthcare applications.
@@ 2.1 Blocks with Spike Neurons
The Transformer-based SNN block, depicted in Figure 3, integrates the attention mechanisms of traditional Transformers with the spiking neurons. In this block, spiking neurons generate spike-based queries $Q_{s}$, keys $K_{s}$, and values $V_{s}$, which are essential components of the self-attention mechanism. These spike-based representations are obtained by processing incoming spike signals through re-parameterized convolution layers that culminate with spiking neurons. Specially, for a input $X$ to the transformer-based SNN block, the queries, keys, and values are computed as
@@ 2.2 Eye-gaze Guided Spike Vision Transformer
We augment the original images by integrating eye-gaze data to highlight key regions without losing information from other areas through the gaze mask (GM) method. Let $\mathbf{I}\in\mathbb{R}^{H\times W\times C}$ denote the original image, and let $\mathbf{M}\in\mathbb{R}^{H\times W}$ be the normalized eye-gaze mask with values between 0 and 1. The enhanced image $\mathbf{I}^{\prime}$ is computed as:
@@
where $\odot$ denotes element-wise multiplication, $\mathbf{1}$ is an all-ones matrix of compatible dimensions, and $\alpha$ is a hyperparameter controlling the emphasis on eye-gaze regions. This operation amplifies pixel intensities in areas with higher eye-gaze values, effectively guiding the network’s focus. To align the model’s attention with human gaze patterns, we introduce an attention alignment (ALH) loss. Let $\mathbf{A}_{t}\in\mathbb{R}^{N\times N}$ be the model’s attention map from the last layer of the final Transformer block, and $\mathbf{A}_{g}\in\mathbb{R}^{N\times N}$ be the attention map derived from eye-gaze data, where $N$ is the number of tokens. The attention alignment loss $\mathcal{L}_{\text{align}}$ is defined as:
@@ 3.2 Results and Analysis
As shown in Table 1, our eye-gaze guided SNN architecture, EG-SpikeFormer achieves state-of-the-art performance across all metrics, outperforming classical ViTs, eye-gaze augmented ViTs, and CNNs. Besides lower energy consumption, our model enhances performance, demonstrating efficiency and superior capability in complex medical imaging tasks. Leveraging the spatio-temporal event-driven computation of SNNs, our architecture balances high performance with low energy consumption without compromising accuracy. It surpasses traditional deep learning architectures, highlighting the potential of SNNs in resource-constrained medical environments. Integrating eye-gaze data effectively guides the model’s attention, improving accuracy and convergence speed without significantly increasing computational resources. We evaluated two training processes: a two-step and a four-step. Even with two-step training, our model matches or exceeds other state-of-the-art methods. The four-step training further enhances capabilities, indicating that additional steps can unlock more potential without substantially increasing energy consumption.
@@ 3.3 Ablation Study
GM ALH T4 Acc. $\uparrow$ AUC $\uparrow$ F1 $\uparrow$ Acc. $\uparrow$ AUC $\uparrow$ F1 $\uparrow$
@@
We conducted an ablation study to assess the impact of GM, ALH, and the four-step training process (T4) on model performance. The detailed results are presented in Table 2. The model without T4 also showed lower performance metrics. The highest performance was achieved when all components were included, confirming that GM, ALH, and T4 each contribute significantly to the model’s effectiveness.
@@ 4 Discussion
Neuromorphic chips, leveraging SNNs as the most accessible approach, offer transformative potential for clinical diagnostics by addressing key challenges such as energy efficiency, data scarcity, and interpretability [25, 26]. EG-SpikeFormer tackles shortcut learning and focuses on diagnostically relevant features, making it ideal for real-time medical applications in resource-constrained settings and offering a forward-thinking framework for medical neuromorphic chip design. Additionally, its eye-gaze-driven, spatio-temporal event processing optimizes neuromorphic systems for clinical tasks. By enhancing diagnostic accuracy and promoting algorithm-hardware co-design, EG-SpikeFormer facilitates the development of efficient, high-performance medical chips. Futhermore, its ability to reduce data demands and computational overhead accelerates energy-efficient and scalable chip design, advancing accessible clinical diagnostics and expanding the role of neuromorphic technologies in healthcare.
@@ 5 Conclusion
EG-SpikeFormer integrates eye-gaze data with SNN to enhance medical image analysis, utilizing the energy efficiency and spatio-temporal dynamics of SNN alongside the feature extraction strengths of CNN and Transformers. By directing attention on clinically relevant regions, the model achieves superior diagnostic accuracy, interpretability, and generalizability, and addresses the limitations of conventional models. These results highlight neuromorphic computing’s potential in healthcare, providing a high-performance, low-energy solution, particularly well-suited for medical environments with critical data constraints and accuracy requirements.